ganeshie8 (ganeshie8):

The wheel shown has eight equally spaced spokes and a radius of 30 cm. It is mounted on a fixed axle and is spinning at 2.5 rev/s. You want to shoot a 20-cm-long arrow parallel to this axle and through the wheel without hitting any of the spokes. Assume that the arrow and the spokes are very thin. (a) What minimum speed must the arrow have? (b) Does it matter where between the axle and rim of the wheel you aim? If so, what is the best location?

ganeshie8 (ganeshie8):



Miracrown (miracrown):

The arrow travels a distance equal to its own length while passing through the wheel, agree? Suppose its speed is V then, the time taken to cross is L/V where L is length of arrow \[t \space = \space \frac{ L }{ V }\]

Miracrown (miracrown):

next, within this time t, the wheel must rotate AT MOST 1/8 th of a full rotation Else, the spoke will hit the arrow somewhere

ganeshie8 (ganeshie8):

so the arrow has t = 20/v seconds to pass through, the best case

Miracrown (miracrown):

Right

Miracrown (miracrown):

Do we know t?

ganeshie8 (ganeshie8):

we have to find it i think

Miracrown (miracrown):

Yep. So, read the question again, you will get it ;)

ganeshie8 (ganeshie8):

Ahh you're so fast ! it seems the wheel can rotate a maximum of 1/8th of revolution in 20/v seconds, so we have \[20/v = \dfrac{\pi/4}{2.5*2\pi} \implies v = 400 cm/sec\]

ganeshie8 (ganeshie8):

ofcourse you're fast, you have an alien head :D

ganeshie8 (ganeshie8):

any ideas on part b

Miracrown (miracrown):

Hahahahahaha

ganeshie8 (ganeshie8):

if it were me with the bow and arrow, my intuition tells me to aim closer to the rim so that i have large space to cover... but my intuition is almost always wrong... so il have to think a bit more...

Miracrown (miracrown):

I will aim to the middle to minimize the accidental hitting on sides

Miracrown (miracrown):

If you are a perfect shooter, it does not matter

Miracrown (miracrown):

any location will be equally good

imqwerty (imqwerty):

:) yes

Miracrown (miracrown):

Just dont aim at the center where there is the axle

ganeshie8 (ganeshie8):

lol we could always aim outside the wheel but no prize for you, thats not what the question is asking haha

ganeshie8 (ganeshie8):

Oh are you saying, it doesn't matter where the arrow is aimed at ?

Miracrown (miracrown):

yes m8

imqwerty (imqwerty):

haha only if ur like best at shooting in wrld :D

Miracrown (miracrown):

it does NOT matter

imqwerty (imqwerty):

it will matter if m shooting :)

Miracrown (miracrown):

me too, i'll prolly shoot right in your eyeball @imqwerty

Miracrown (miracrown):

oh, oooops

ganeshie8 (ganeshie8):

may i know why it wouldn't matter

imqwerty (imqwerty):

ahahaha cx

Miracrown (miracrown):

you're one curious man ganeash

ganeshie8 (ganeshie8):

i know you're not kidding because the answer on the back of textbook says that it doesn't matter

imqwerty (imqwerty):

:O

Miracrown (miracrown):

but let's be real, b/c the location does not enter our equations The V we calculated does not depend on the radial location

Miracrown (miracrown):

Another way to look at it is this: As you go farther out along the radius, the rotational velocity increases. However, the available distance along the circumference also increases PROPORTIONATELY.

Miracrown (miracrown):

This basically render the radial position irrelevant

ganeshie8 (ganeshie8):

Wow! I completely missed that! angular velocity is 2.5 rev/s, which is constant, so all the particles on a spoke move by same angle in a given time

imqwerty (imqwerty):

yes thas correct explanation :) what i said was wrng the time taken was equal but not the velocity

Miracrown (miracrown):

;)

hartnn (hartnn):

This alien is a genius! making full use of her huge head :P

imqwerty (imqwerty):

tru that

Miracrown (miracrown):

I try ma best :P

ganeshie8 (ganeshie8):

it's power comes from extracting energy from human brains

imqwerty (imqwerty):

LOL

hartnn (hartnn):

O.o *stays away from that alien then*

Miracrown (miracrown):

even you can use your brain to its full potential, there is just a small twist which I ain't revealing, uh-nuh

OpenStudy (mom.):

(:

OpenStudy (kainui):

I guess I have to unfan @Miracrown so that I can fan her now.

ganeshie8 (ganeshie8):

thinking a bit more, below is so wrong : ``` Wow! I completely missed that! angular velocity is 2.5 rev/s, which is constant, so all the particles on a spoke move by same angle in a given time ``` i think the actual reason that the placement doesn't matter is the fact that it is a "rigid body". Even if the angular velocity changes, it changes the same for all the particles on the spoke. So constant angular velocity has nothing to do with it.

OpenStudy (mrnood):

It is evident that all points on 1 spoke rotate to th epositin of the previous spoke in the same time - other wise the wheel would disintegrate

ganeshie8 (ganeshie8):

Exactly!